# A Goal-Oriented Reinforcement Learning-Based Path Planning Algorithm for Modular Self-Reconfigurable Satellites

### v1

@@ 1 Introduction
Compared to the traditional single-satellite single-mission paradigm, MSRSs offer notable advantages: they streamline satellite design and manufacturing through standardized modules, enhance deployment efficiency of space systems such as space telescopes (Battista and Helvajian, 2020) and orbital solar power stations (Helvajian, 2020) via batch launches and autonomous on-orbit assembly, and extend operational lifespan of space assets by enabling in-orbit repair, refueling, and recovery (Shi et al., 2022; Wang et al., 2023; Liu et al., 2020; Jiang et al., 2022) , thereby mitigating space debris accumulation.
@@ 3.3.2 Reward Function Design
To address the aforementioned issue, we carefully designed the following reward function:
@@
In the above function, $\mathcal{S}$ and $\mathcal{S^{\prime}}$ represent the state space of the configuration before and after the action, respectively. With this design, the agent can effectively evaluate whether its current action moves the module closer to the target configuration (positive reward) or further away (negative reward). In addition, we assign a reward of -1 when the agent chooses a ”no-op” (no movement) action before reaching the target. This discourages the agent from falling into suboptimal strategies where it remains stationary without completing the reconfiguration.
@@ 4.1 Soft Actor-Critic
Soft Actor-Critic (SAC) (Haarnoja et al., 2018) is a deep reinforcement learning algorithm based on the maximum entropy framework. It adopts an actor-critic architecture to maximize cumulative reward while encouraging exploration by maximizing policy entropy, helping prevent premature convergence to suboptimal solutions. Compared to the standard objective in Equation 1, the cumulative reward that SAC aims to maximize is modified as follows:
@@ 4.2 SAC for Discrete Action in GORL
         Update critic networks using Eq. (7)
@@
              Update actor network using Eq. (9)
              Update $\alpha$ using Eq. (10)
@@ 5 Experiments
The experimental results demonstrate that, compared to models without HER or action masking, our method is able to identify optimal strategies more efficiently and avoids policy entropy divergence caused by convergence to suboptimal strategies. Figure 7 illustrates an example of the reconfiguration produced by our algorithm.
@@ 6 Conclusion
The proposed model is capable of solving the path planning problem for any target configuration of electromagnetic rotating MSRSs, effectively addressing the generalization limitation of previous studies. However, the current approach still suffers from high training time costs. Future work may explore two potential directions for improvement: (1) utilizing $A^{*}$ algorithm to pre-sample trajectory , and (2) considering applying distributed action modeling methods to reduce the dimensionality of the action space.
@@
the National Natural Science Foundation of China(62203145), National Key R&D Program of China(2021YFC2202900), the China Postdoctoral Science Foundation(2022M710948), the Civil Aerospace Research Project of China(D030312).

### v2

@@ 1 Introduction
Compared to the traditional single-satellite single-mission paradigm, MSRSs offer notable advantages: they streamline satellite design and manufacturing through standardized modules, enhance deployment efficiency of space systems such as space telescopes (Battista and Helvajian, 2020) and orbital solar power stations (Helvajian, 2020) via batch launches and autonomous on-orbit assembly, and extend operational lifespan of space assets by enabling in-orbit repair, refueling, and recovery (Shi et al., 2022; Wang et al., 2023; Jiang et al., 2022) , thereby mitigating space debris accumulation.
@@ 3.3.2 Reward Function Design
To address the aforementioned issue, we carefully designed the following reward function through empirical analysis:
@@
In the above function, $\mathcal{S}$ and $\mathcal{S^{\prime}}$ represent the state space of the configuration before and after the action, respectively. With this design, the agent can effectively evaluate whether its current action moves the module closer to the target configuration (positive reward) or further away (negative reward). In addition, we assign a reward of -1 when the agent chooses a “no-op” (no movement) action before reaching the target. This discourages the agent from falling into suboptimal strategies where it remains stationary without completing the reconfiguration.
@@ 4.1 Soft Actor-Critic
Soft Actor-Critic (SAC) (Haarnoja et al., 2018) is a deep reinforcement learning algorithm based on the maximum entropy framework. It adopts an actor-critic architecture to maximize cumulative reward while encouraging exploration by maximizing policy entropy, helping prevent premature convergence to suboptimal solutions. Compared to the standard objective in (1), the cumulative reward that SAC aims to maximize is modified as follows:
@@ 4.2 SAC for Discrete Action in GORL
         Update critic networks using (7)
@@
              Update actor network using (9)
              Update $\alpha$ using (10)
@@ 5 Experiments
The experimental results demonstrate that, compared to models without HER or action masking, our method is able to identify optimal strategies more efficiently and avoids policy entropy divergence caused by convergence to suboptimal strategies. Figure 7 illustrates an example of the reconfiguration produced by our algorithm. The code of the proposed approach and the learned models are available at https://github.com/perfactliu/GORL4MSRS.
@@ 6 Conclusion
The proposed model is capable of solving the path planning problem for any target configuration of electromagnetic rotating MSRSs, effectively addressing the generalization limitation of previous studies. However, the current approach still suffers from high training time costs, especially when tackling larger modular satellite systems. Future work may explore two potential directions for improvement: (1) utilizing $A^{*}$ algorithm to pre-sample trajectory , and (2) considering applying distributed action modeling methods to reduce the dimensionality of the action space.
@@
National Key R&D Program of China(2021YFC2202900), the National Natural Science Foundation of China(62203145), the China Postdoctoral Science Foundation(2022M710948), the Civil Aerospace Research Project of China(D030312).